an experience bricklayer requires 20 h to complete a job. after the bricklayer and an apprentice work for 12 h, the bricklayer moves to another job. the apprentice requires 12 more hours to finish the job. how long would it have taken the apprentice working alone to do the job?
Answer by rosey51235(30)

Hi :)
I probably will explain this weirdly, so I'll try but I'll start with an apology. Sorry:)
Anyways, if the bricklayer needs 20hrs and he stopped after 12, thats 12/20=6/10=3/5 of the job.
That means there's 2/5 of the job left, and 12hrs. divided by 2( TWO fifths) is 6.
If 6 is only 1/5 of the time it took the apprentice, then alone (needing to do 5/5 of the job) you multiply 6 by 5.
6*5=30, so alone, it would've taken the apprentice 30 hours to finish the job.
Hope that wasn't too hard to follow, because even though I get problems right, my math teachers tell me I always manage to solve it in a way that WILL work, but isn't the usual route. :p
Hope I helped! :)
-Rose